OpenStudy (angelwings996):

A toy company is considering a cube or sphere-shaped container for packaging a new product. The height of the cube would equal the diameter of the sphere. Compare the volume-to-surface area ratios of the containers. Which packaging will be more efficient?

OpenStudy (angelwings996):

For a sphere \[SA = 4 \pi r ^{2}\]

OpenStudy (campbell_st):

ok... so the volume of the cube is \[V = x^3 \] for the shpere if the diameter is x then the radius is x/2 the volume of the sphere is \[V = \frac{4}{3} \pi r^3\] substitute r = x/2 and evaluate

OpenStudy (campbell_st):

surface area of the cube is \[SA = 6x^2\] for the surface are of the sphere its \[SA = 4 \pi r^2\] substitute r = x/2 hope this helps..

OpenStudy (angelwings996):

@campbell_st so would I use both the surface area formulas and the volume formulas for this ?

OpenStudy (campbell_st):

they are asking you to find a ratio Volume to Surface area so to me I do Cube \[\frac{V}{SA} = \frac{x^3}{6x^2} = \frac{x}{6}\] you would need a similar relationship for the sphere... then I'd assume the smaller number would be more efficient...

OpenStudy (angelwings996):

okay so would the sphere one look like \[\frac{ V }{ SA } = \frac{ 4/3 }{ 4 } = 3 ?\]

OpenStudy (angelwings996):

Im kind of confused on how you would do the pi and radius sign in here

OpenStudy (campbell_st):

yep... so volume of the sphere is \[V = \frac{4}{3} \pi (\frac{x}{2})^3 = \frac{\pi x^3}{6}\] and \[SA = \pi x^2\] the interesting thing is the ratios are equal... Sphere.. \[\frac{V}{SA} = \frac{\frac{\pi x^3}{6}}{\pi x^2}\]

OpenStudy (angelwings996):

Ohh so the answer would be that they are both equally efficient?

OpenStudy (campbell_st):

well thats what the math says... providing there are correct... the ratio for the Sphere is \[\frac{V}{SA} = \frac{x}{6}\]

OpenStudy (campbell_st):

oops should be providing the math is correct...

OpenStudy (angelwings996):

Ohh okay! Well thank you so much for your help..again lol

OpenStudy (campbell_st):

just hope it didn't confuse... good luck

OpenStudy (angelwings996):

No, I understand, a lot better than before actually so thank you